OpenStudy (anonymous):

9x + 2y = 6 in slope intercept form

OpenStudy (anonymous):

Isolate the y variable

OpenStudy (anonymous):

What does isolate mean?

OpenStudy (anonymous):

9x + 2y = 6 2y = -9x + 6 y = -9/2x + 3

OpenStudy (anonymous):

what grade are you in?

OpenStudy (anonymous):

10th grade i know how to put stuff in standard form but I get confused with the slope intercept form

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Alright thank you